# Enhancing Sequential Recommendations through Multi-Perspective Reflections and Iteration

### v1

@@ 3.5. Reflection Selection
HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10 HR@5 HR@10 N@5 N@10
Traditional Deep Recommendations Methods
Caser 0.2555 0.3603 0.1823 0.2161 0.3024 0.4397 0.1997 0.2440 0.2931 0.3968 0.2193 0.2527
FDSA 0.2536 0.3670 0.1875 0.2241 0.3073 0.4466 0.2150 0.2597 0.2842 0.3905 0.2157 0.2497
BERT4Rec 0.2402 0.3613 0.1680 0.2072 0.2796 0.4160 0.1861 0.2299 0.2618 0.3789 0.1954 0.2326
GRU4Rec 0.2431 0.3708 0.1722 0.2129 0.3192 0.4644 0.2083 0.2554 0.3021 0.4021 0.2201 0.2520
SASRec 0.3570 0.4280 0.2788 0.3016 0.2915 0.4348 0.1898 0.2369 0.3450 0.3890 0.2755 0.2896
Training-Free LLM-based Methods
LLM4RS 0.1087 0.1277 0.0826 0.0887 0.1976 0.2233 0.1338 0.1421 0.0900 0.1188 0.0624 0.0717
LLMRank(CoT) 0.3337 0.4061 0.2595 0.2828 0.3508 0.4496 0.2607 0.2925 0.2904 0.3709 0.2202 0.2461
Training-Required LLM-based Methods
LC-Rec 0.3768 0.4309 0.3006 0.3181 0.4140 0.4646 0.3426 0.3593 0.3296 0.3920 0.2486 0.2691
BinLLM 0.2422 0.2896 0.1786 0.2126 0.2701 0.3087 0.2119 0.2523 0.2386 0.2799 0.1958 0.2442
Re2LLM 0.4252 0.4938 0.3535 0.3757 0.4180 0.5198 0.3231 0.3559 0.3521 0.4424 0.2787 0.3074
MoRE 0.4376 0.5043 0.3705 0.3922 0.4773 0.5731 0.3647 0.3957 0.3824 0.4674 0.2981 0.3251
Imp. 2.92% 2.13% 4.81% 4.39% 14.19% 10.25% 6.45% 10.13% 8.61% 5.65% 6.96% 5.76%
Table 1. Recommendation performance comparison. The best and the second-best performances are denoted in bold and underlined fonts, respectively. “N@K” is short for “NDCG@K”. The “Imp.” indicates the percentage improvement of MoRE over the best performances from baselines. denotes MoRE performs significantly better than baselines based on two-tailed paired t-test with Bonferroni correction ($p<0.05$).
4. Experiment
In this section, we conduct extensive experiments to answer the following research questions:
  • RQ1: How does MoRE compare to various existing baselines in terms of recommendation performance?
  • RQ2: How effective are the multi-perspective reflections and reflection selection in MoRE?
  • RQ3: For reflection memory maintenance, how effective are the refining and iteration strategies?
  • RQ4: Does MoRE have an advantage in terms of training time overhead and GPU memory usage?
@@ 4.1. Experiment Settings
Dataset. We conduct our experiments on three product categories from the Amazon dataset (Ni et al., 2019): “Arts”, “Video”, and “Instruments”. Each item in these categories has attributes that reflect both explicit user preferences (e.g., titles) and implicit preferences (e.g., brands). To provide sequential recommendations, we follow the approach in (Zheng et al., 2024) and filter out unpopular users and items with fewer than five interactions. The statistics of the pre-processed datasets can be found in App. A.1.
Baselines. We adopt three categories of baselines: traditional deep recommendation methods (e.g., SASRec (Kang and McAuley, 2018), etc.), training-free LLM-based methods (e.g., LLM4RS (Dai et al., 2023), etc.), and training-required LLM-based methods (e.g., Re2LLM (Wang et al., 2024), etc.). See App. A.2 for details.
@@
Evaluation Metrics. To evaluate the performance of our MoRE and other baselines, we adopt two widely used metrics, top-$k$ Hit Ratio (HR@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$), with $k\in\{5,10\}$. Following (Zheng et al., 2024; Zhou et al., 2022), we employ the leave-one-out strategy for the obtaining of training, validation, and test data. Specifically, for each user behavior sequence, the last item is used as the test data, the penultimate item is used as the validation data, and the remaining interaction records are used for training. Following previous work (Wang et al., 2024), we set every user’s candidate set size $|\mathcal{C}_{u}|$ to 50, where we sample 49 negative items (i.e. items which users do not interact with) from the whole item set for each positive item (i.e. the target item). We perform a ranking evaluation over the candidate set. MoRE and all baselines are required to give a top-10 ranked list for each user.
Implementation Details. We use LLaMa-3111https://huggingface.co/gradientai/Meta-Llama-3-8B-bf16.  (Dubey et al., 2024) as the LLM backbone for MoRE and LLM-based baselines. Implementation details can be found in App. A.3.
4.2. Recommendation Performance
To answer RQ1, we compare the SeqRec performance of MoRE against baselines for 10 trials, with the average results presented in Tab. 1. We can observe the following:
(1) MoRE outperforms all other baselines across each metric on each dataset. On all metrics across three datasets, MoRE outperforms the best baseline by average improvements of 3.56%, 10.25%, and 6.75%, respectively. We attribute this to our proposed LLM-based reflection framework which leverages the reflection capabilities of LLMs to understand both users’ explicit and implicit preferences. Additionally, it enables LLMs to grasp the CF knowledge in a straightforward and comprehensible manner, leading to improved performance in SeqRec.
(2) In general, LLM-based methods demonstrate performance comparable to the traditional methods, which validates the potential of LLMs in SeqRec. Additionally, LLMRank with CoT (Feng et al., 2024) significantly outperforms LLM4RS, indicating that prompt engineering strategies like CoT (Wei et al., 2022) can substantially enhance performance without fine-tuning.
(3) Typically, trained methods demonstrate superior performance compared to their untrained counterparts. Surprisingly, BinLLM delivers relatively underwhelming outcomes. Designed to leverage binary sequences in prompts to capture the correlation between a user-item pair, BinLLM struggles with the nature of SeqRec, which involves a user and multiple candidate items. The insertion of such sequences leads to excessively long prompts, hindering LLMs’ capacity to accurately identify user-item associations.
In summary, MoRE exhibits superior performance in SeqRec, thereby validating the efficacy of our approach.
4.3. Analysis of Multi-Perspective Reflection and Selection
To validate the effectiveness of each perspective of reflection and answer RQ2, we conduct an analytical experiment on Amazon Arts. The results are presented in Tab. 3.
We find that MoRE, equipped with the reflection selection, outperforms alternatives, demonstrating the effectiveness of the multi-perspective reflections and the corresponding perspective selection. Specifically, MoRE outperforms the variants with not only the single reflection perspective but also the dual or triple perspectives simultaneously. This underscores that no single perspective of reflection is universally optimal for all users, thereby confirming the necessity for selective choice among perspectives.
Moreover, MoRE’s selection strategy surpasses both the Random and Greedy selection approaches which respectively denote a random reflection selection from the three memory banks and the selection solely based on downstream recommendation performance. This highlights the effectiveness of our contextual bandit modeling, enabling MoRE to judiciously choose the most suitable reflection for each user.
We also conduct a case study in App. A.5 to further assess the reflection selection of MoRE in comparison to other alternative selection strategies.
@@ 4.3. Analysis of Multi-Perspective Reflection and Selection
Figure 3. Impact of refining and iteration strategies on recommendation performance. We repeated the experiment 5 times, with the mean value for each method represented by the line and the shaded area indicating the standard deviation.
Methods Arts Games Instruments
Time/h GRAM/GB Time/h GRAM/GB Time/h GRAM/GB
BinLLM (Fine-Tune) 10.209(±0.090) 124.67 11.468(±0.072) 120.57 10.435(±0.094) 115.48
Re2LLM (RL) 4.007(±0.007) 2.77 3.894(±0.005) 1.99 4.152(±0.008) 1.13
MoRE (RL) 3.982(±0.001) 2.77 3.854(±0.002) 1.99 4.093(±0.003) 1.13
Table 2. Training cost comparisons among LLM-based methods on four A800 GPUs. ”Time/h” represents the training time, shown as the mean and standard deviation (SD). ”GRAM/GB” indicates the maximum memory consumption.
Methods N@5 N@10
MoRE 0.3705 0.3922
- Random 0.3421 0.3701
- Greedy 0.3528 0.3786
- $Ref_{\mathrm{EP}}$ only 0.3670 0.3881
- $Ref_{\mathrm{IP}}$ only 0.3571 0.3788
- $Ref_{\mathrm{CF}}$ only 0.3572 0.3788
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}$ 0.3443 0.3651
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{CF}}$ 0.3311 0.3533
- $Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3033 0.3217
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.3051 0.3268
Table 3. Ablation studies on Amazon Arts. “Random” indicates a random reflection selection from the three memory banks, while “Greedy” denotes that reflection selection is based solely on downstream SeqRec on the validation set.
4.4. Analysis of the Refining and Iteration
To answer RQ3, we test the impact of three refining strategies on SeqRec under different iteration rounds, taking reflections on explicit user preferences as an example. From the Fig. 3, we observed that:
@@ 4.4. Analysis of the Refining and Iteration
(2) Refining: Refining at group and individual levels shows a significantly better trend over that at the global level during iterations, particularly on Arts (Fig. 3(a)). This improvement may be due to the global-level reflections introducing noise from other users, whereas group- and individual-level demonstrations more effectively elicit personalized reflective capabilities from the reflectors.
Additionally, we found that the refining and iteration strategies are most effective for reflections on explicit preferences, compared to reflections on implicit preferences and collaborative filtering. Detailed experiments and discussions are presented in App. A.4.
4.5. Analysis of LLM Training Cost
To address RQ4, we conduct 5 repeat experiments to compare the training time and maximum memory consumption associated with all the LLM-based baselines that require training. To mitigate the effects of experimental randomness, we perform 5 repetitions of the experiments and report the results as the mean and standard deviation (SD), as shown in Tab. 2. Our findings are as follows:
(1) The training cost of reflection-based methods is lower than fine-tuning methods. This is because the reflection-based methods do not require fine-tuning, thus avoiding significant training time and memory costs.
(2) Under reflection-based setup, MoRE has lower training time and memory consumption compared to Re2LLM. We attribute this to the larger action space of Re2LLM. Specifically, MoRE’s Meta-Reflector does not require selecting specific reflections but only needs to choose among three directions of reflection. In contrast, Re2LLM must first acquire a knowledge base of reflection and, once fixed, conduct PPO training of the retrieval agent within that knowledge base. This means it selects specific reflections from the entire knowledge base, with a larger action space.
In summary, MoRE has the lowest training costs including time overhead and GPU memory usage among all the LLM-based baselines, demonstrating its training efficiency.

### v2

@@ 4. Experiment
TALLRec 0.2559 0.1975 0.1588
BinLLM 0.2401 0.2347 0.1875
LLaRA 0.2803 0.2263 0.1763
A-LLMREC 0.2193 0.2342 0.1897
MoRE 0.2898 0.2569 0.2163
Table 4. Ablation studies on Amazon Arts. “Random” indicates a random reflection selection from the three memory banks, while “Greedy” denotes that reflection selection is based solely on downstream SeqRec on the validation set. $Ref_{\mathrm{all}}$ denotes $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$. Note: “+” denotes concatenation, which may cause conflicting reflections leading to suboptimal performance. MoRE achieves the best results through user-personalized RL selection.
Methods HR@1 HR@5 HR@10 N@5 N@10
MoRE 0.2898 0.4376 0.5043 0.3705 0.3922
- Random 0.2736 0.4023 0.4881 0.3421 0.3701
- Greedy 0.2812 0.4156 0.4957 0.3528 0.3786
- $Ref_{\mathrm{EP}}$ only 0.2850 0.4356 0.5033 0.3670 0.3881
- $Ref_{\mathrm{IP}}$ only 0.2850 0.4214 0.4881 0.3571 0.3788
- $Ref_{\mathrm{CF}}$ only 0.2850 0.4290 0.4871 0.3572 0.3788
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{IP}}$ 0.2583 0.4071 0.4700 0.3443 0.3651
- $Ref_{\mathrm{EP}}+Ref_{\mathrm{CF}}$ 0.2707 0.3927 0.4623 0.3311 0.3533
- $Ref_{\mathrm{IP}}+Ref_{\mathrm{CF}}$ 0.2402 0.3584 0.4156 0.3033 0.3217
- $Ref_{\mathrm{all}}$ 0.2393 0.3594 0.4271 0.3051 0.3268
@@ 4.1. Experiment Settings
Dataset. Following (Zheng et al., 2024), we conduct our experiments on three subsets of the Amazon dataset (Ni et al., 2019): “Arts”, “Video”, and “Instruments” [^1] Note that “Arts”, “Games” and “Instruments” represent the abbreviations of “Arts, Crafts and Sewing”, “Video Games” and “Musical Instruments” respectively in the Amazon Review dataset (2018).. Each item in these subsets contains attributes that capture both explicit user preferences (e.g., item titles) and implicit preferences (e.g., attributes like brands and functional traits ). We follow (Zheng et al., 2024) to preprocess the data to achieve a fair comparison, with the statistical information shown in Tab. 1.
Baselines. We adopt three types of SeqRec methods as baselines: (1) Traditional Deep Learning Models: Caser (Tang and Wang, 2018) captures local and global patterns via CNN; GRU4Rec (Hidasi et al., 2015) models sequences with GRUs; SASRec (Kang and McAuley, 2018) and BERT4Rec (Sun et al., 2019) employ uni-/bidirectional Transformers for next-item prediction; FDSA (Zhang et al., 2019) models item-feature transitions via dual attention. (2) Training-Free LLM Methods: LLM4RS (Dai et al., 2023) aligns LLMs with IR ranking strategies; LLMRank (Hou et al., 2024) achieves zero-shot ranking via specialized prompting. (3) Training-Required LLM Methods: LC-Rec (Zheng et al., 2024) expands vocabulary for collaborative filtering; TALLRec (Bao et al., 2023) uses instruction tuning with text descriptions; BinLLM (Zhang et al., 2024a) encodes user-item interactions as binary sequences; Re2LLM (Wang et al., 2024) employs fixed reflection pools for suggestions; LLaRA (Liao et al., 2024) combines user knowledge and behavior patterns via LoRA; A-LLMREC (Kim et al., 2024) aligns CF embeddings with LLM vocabulary.
@@
Evaluation Metrics. We adopt two widely used metrics, top-$k$ Hit Ratio (HR@$k$) and Normalized Discounted Cumulative Gain (NDCG@$k$), with $k\in\{5,10\}$. We follow (Zheng et al., 2024; Zhou et al., 2022) to employ the leave-one-out strategy for the obtaining of training, validation, and test data. Specifically, for each user behavior sequence, the last item is used as the test data, the penultimate item is used as the validation data, and the remaining interaction records are used for training. Following (Hou et al., 2024; Wang et al., 2024; Kim et al., 2024; Liao et al., 2024), we randomly sample $|\mathcal{C}_{u}|-1$ negative items to construct $|\mathcal{C}_{u}|$ candidates with 1 target item for each user. Following  (Wang et al., 2024), $|\mathcal{C}_{u}|$ is set to 50 for all methods. All methods except those inherently limited to top-1 prediction (without ranking) must provide top-10 ranked items.
[…]
Table 5. Training cost comparisons among LLM-based methods on four A800. “Training time” is measured in hours, shown as the mean and standard deviation (SD). “VRAM” denotes the Peak VRAM Usage, measured in GB. MoRE’s RL framework minimizes computation cost through compact action space ($|\mathcal{A}|=3$, 3 perspectives).
Methods Arts Games Instruments
Training time VRAM Training time VRAM Training time VRAM
TALLRec (Fine-Tune) 26.228(±0.818) 46.98 18.018(±0.290) 43.24 18.338(±0.487) 42.95
BinLLM (Fine-Tune) 10.209(±0.540) 124.67 11.468(±0.328) 120.57 10.435(±0.451) 115.48
LLaRA (Fine-Tune) 9.013(±0.161) 77.92 6.405(±0.281) 65.83 6.863(±0.187) 74.36
A-LLMREC (Fine-Tune) 56.975(±1.621) 217.74 55.975(±1.058) 202.09 56.316(±1.992) 231.63
Re2LLM (RL, $|\mathcal{A}|=20$) 4.007(±0.007) 2.77 3.894(±0.005) 1.99 4.152(±0.008) 1.13
MoRE (RL, $|\mathcal{A}|=3$) 3.982(±0.001) 2.77 3.854(±0.002) 1.99 4.093(±0.003) 1.13
@@
(d) Iteration for 3 Perspectives
Figure 4. Impact of refining and iteration strategies on recommendation performance. We repeated the experiment 5 times, with the mean value for each method represented by the line and the shaded area indicating the standard deviation. (a), (b), and (c) demonstrate the effects of iterations and different levels of refining on three datasets, while (d) shows the effects of the three perspectives with Games as an example.
4.2. Recommendation Performance Comparison
To answer RQ1, we compare the SeqRec performance of MoRE against baselines for 10 trials, with the average results presented in Tab. 2 and Tab. 3. We have the following observations: (1) MoRE outperforms all baselines across all metrics on all datasets, except for the suboptimal performance on the HR@1 on “Games”. (2) In general, LLM-based methods demonstrate performance comparable to the traditional methods, which validates the potential of LLMs in SeqRec. (3) Notably, MoRE outperforms LLM-based methods that specialize in predicting one single item (HR@1 only), even when those methods sacrifice ranking capabilities444 Note that these methods are inherently limited to top-1 predictions. Therefore, for these methods, we retain only top-1 results for a fair comparison here..
In summary, MoRE exhibits superior performance in SeqRec, validating the efficacy of our approach. We attribute this to our proposed LLM-based reflection framework, which leverages LLM’s analytical and reasoning capabilities to capture users’ explicit preferences and implicit behavioral patterns while seamlessly integrating CF knowledge in a clear and comprehensible manner, ultimately enhancing performance in sequence recommendation tasks.
Figure 5. Example of how LLM reflections from explicit preference (EP), implicit preference (IP), and collaborative filtering (CF) perspectives improve recommendation.
4.3. Ablation Study: Multi-Perspective Reflection and Selection
To validate the effectiveness of each perspective of reflection and answer RQ2, we conduct an ablation study on Amazon Arts. The results are presented in Tab. 4.
We find that MoRE, equipped with the reflection perspective selection, outperforms alternatives, demonstrating the effectiveness of the multi-perspective reflections and the corresponding perspective selection. Specifically, MoRE outperforms the variants with not only the single reflection perspective but also the dual or triple perspectives simultaneously. The simultaneous use of multiple perspectives leads to a decrease in performance. This implies potential discrepancies among perspectives and no single perspective of reflection is universally optimal for all users, thereby confirming the necessity for selective choice among perspectives.
Moreover, MoRE’s selection strategy surpasses both the Random and Greedy selection approaches which respectively denote a random reflection selection from the three memory banks and the selection solely based on downstream recommendation performance. This highlights the effectiveness of our contextual bandit modeling, enabling MoRE to choose the most suitable reflection for each user.
4.4. Ablation Study: Refining and Iteration
To answer RQ3, we test the impact of three refining strategies on SeqRec under different iteration rounds, taking reflections on explicit features as an example. From Fig. 4, we find:
@@ 4.4. Ablation Study: Refining and Iteration
(2) Refining: Refining at group and individual levels shows a significantly better trend over that at the global level during iterations, particularly on Arts (Fig. 4(a)). We assume that the global-level reflections may introduce noise from other users, whereas group- and individual-level demonstrations more effectively elicit personalized reflective capabilities from the reflectors.
4.5. Analysis of Computational Cost
To address RQ4, we compare training costs (time and GPU memory) across LLM-based baselines using five repeated experiments, reported as Avg. $\pm$ SD [^5] Average and standard deviation. in Tab. 5. We find: 1) Reflection-based methods (e.g., MoRE) incur lower costs than fine-tuning approaches by avoiding LLM parameter updates; 2) MoRE outperforms Re2LLM in efficiency due to its smaller action space. MoRE’s Meta-Reflector selects from 3 reflection perspectives, while Re2LLM trains a PPO-based retrieval agent over a fixed base [^6] The size is 20., requiring larger action-space exploration. In summary, MoRE achieves the lowest training costs (time and memory) among all baselines, demonstrating superior training efficiency.
4.6. Case Study
We present a case study demonstrating reflections from EP, IP, and CF perspectives to enhance LLMREC (Fig. 5). The three perspectives improve performance as follows: EP Reflection focuses on recent interaction trends (e.g., game genre preferences) to align with explicit preferences. This results in the target item being recommended at the 8-th position (e.g., Street Fighter 5), effectively capturing short-term user intentions. IP Reflection identifies patterns in historical attributes (e.g., multiplayer functionality, ”PlayStation” brand) to adapt to implicit preferences. This moves the target to the 6-th position by analyzing deeper user behavioral patterns. CF Reflection leverages collaborative filtering signals and emphasizes highly-rated items, reducing diversity while incorporating historical ratings and relevant sequels. This achieves the top-1 recommendation position. The Meta-reflector dynamically selects the optimal reflection (CF in this case), enabling MoRE to deliver optimal recommendations.
The reflection mechanism bridges the semantic gap between LLMs and CF: LLMs utilize simple score comparisons instead of interpreting embeddings. This approach cost-effectively equips LLMs with domain-specific knowledge (e.g., CF signals) while mitigating their adaptation challenges in specialized domains.